OpenStudy (anonymous):

what is the answer for x^3y^3-a^3 from binomial to trinomial? i need your help now :)

OpenStudy (anonymous):

x^3y^3-a^3=(xy)^3-a^3 and try with this \[A^3-B^3=(A-B)(A^2+AB+B^2)\]